OpenStudy (mollynicole2015):

Solve for z: z 2 = ------ 3

OpenStudy (anonymous):

z = 6

Nnesha (nnesha):

how would you get rid of the fraction ?

OpenStudy (mollynicole2015):

Multiplication?

Nnesha (nnesha):

right do it :=))

OpenStudy (anonymous):

z= 6

OpenStudy (anonymous):

You multiply the denominator by the whole number to get the numerator.

OpenStudy (mollynicole2015):

Okay, so 3 x 3 cancels out, and 2 x 3 is 6 so z = 6.

OpenStudy (youngstudier):

Yep.

OpenStudy (anonymous):

Molly i know a maath wbsite that will give you evey answer for every problem

OpenStudy (mollynicole2015):

What is it. O_O

Nnesha (nnesha):

3 times 3 can't cancels out it's \[\frac{ z }{ 3} \times 3\] which is same as \[z\frac{1}{3} \times 3\] we are multiplying by 3 to cancel out 3 from the denominator

Nnesha (nnesha):

\[z\frac{1}{\cancel{3}} \times \cancel{3}\] 3/3 cancels out

OpenStudy (anonymous):

check inbox